OpenStudy (anonymous):

A student wants to make a pendulum that makes one complete swing, away and back, every 1.5 s. What should the student choose for the mass of the bob and the length of the string?

OpenStudy (anonymous):

Period should be independent of mass. Period is equal to 2 pi x root (l/g) so if T = 1.5 and g = 9.81 l = (1.5 / 2, pi) ^2 9.81 = 0.559m

OpenStudy (anonymous):

Thank you!! =)